In a Gallop Poll of 1487 adults 639 of them had Facebook pages. Find the best point estimate of the percentage of all adults to the nearest 1% who have a Facebook Page.
Answer #1

solution:

Solution :

Given that,

n = 1487

x = 639

Point estimate = = x / n = 639/1487=0.430

which is 43%=0.430